Let A = (-1 0 4)
6 3 2
2 5 6 3x3 then find
A/ Det(A)
B/ Cofactor of A
C/ adjoint of A
D/ Inverse of A
1
Expert's answer
2020-01-30T12:56:09-0500

We have matrix A = "\left(\begin{matrix}-1 & 0 & 4\\6 & 3 & 2\\2 & 5 & 6\end{matrix}\right)";

Solution:

A) We compute determinant of A by using Laplace expansion (in this case for 3x3 matrix) "Det(A) = \sum_{j=1}^3 (-1)^{1+j}a_{1j}M_{1j}," where "M_{1j}" is a minor of A.

So,

"Det(A) = -1 * \left[\begin{matrix}3 & 2\\5 & 6\end{matrix}\right] - 0 * \left[\begin{matrix}6 & 2\\2 & 6\end{matrix}\right] + 4*\left[\begin{matrix}6 & 3\\2 & 5\end{matrix}\right]" "= -1*(18 - 10)+4*(30-6) = 88"

As a result

"Det(A) = 88".

B) Let's call cofactor matrix of A as B. The B can be computed with next expression "b_{ij} = (-1)^{i+j}a_{ij}M_{ij}." Following the given formula, we obtain: "b_{11} = (-1)^{1+1} \left[\begin{matrix}3 & 2\\5 & 6\end{matrix}\right] = 8;" "b_{12} = (-1)^{1+2} \left[\begin{matrix}6 & 2\\2 & 6\end{matrix}\right] =-32;"

"b_{13} = (-1)^{1+3} \left[\begin{matrix}6 & 3\\2 & 5\end{matrix}\right] =24;" "b_{21} = (-1)^{2+1} \left[\begin{matrix}0 & 4\\5 & 6\end{matrix}\right] =20;"

"b_{22} = (-1)^{2+2} \left[\begin{matrix}-1 & 4\\2 & 6\end{matrix}\right] =-14;" "b_{23} = (-1)^{2+3} \left[\begin{matrix}-1 & 0\\2 & 5\end{matrix}\right] = 5;"

"b_{31} = (-1)^{3+1} \left[\begin{matrix}0 & 4\\3 & 2\end{matrix}\right] =-12;" "b_{32} = (-1)^{3+2} \left[\begin{matrix}-1 & 4\\6 & 2\end{matrix}\right] =26;"

"b_{33} = (-1)^{3+3} \left[\begin{matrix}-1 & 0\\6 & 3\end{matrix}\right] =-3;"

Finally, we have

"B = \left(\begin{matrix} 8 & -32 & 24\\20 & -14 & 5\\-12 & 26 & -3\end{matrix}\right)".

C) By definition adjoint (also called "adjugate") is just a transpose matrix of cofactor matrix A

"adj(A) = B^T."

In this case

"adj(A) = \left(\begin{matrix} 8 & 20 & -12\\-32 & -14 & 26\\24 & 5 & -3\end{matrix}\right)."

D) Using the expression "A^{-1} = \frac{1}{det(A)}adj(A)", we obtain

"A^{-1}=\frac{1}{88}\left(\begin{matrix} 8 & 20 & -12\\-32 & -14 & 26\\24 & 5 & -3\end{matrix}\right) =" "\left[\begin{matrix}\frac{1}{11} & \frac{5}{22} & - \frac{3}{22}\\- \frac{4}{11} & - \frac{7}{44} & \frac{13}{44}\\\frac{3}{11} & \frac{5}{88} & - \frac{3}{88}\end{matrix}\right]"